OpenStudy (anonymous):

solve the following logarithmic equation log(6x+2)=1 + log (x-5) find the domian of the composite function f.g f(x) x^2 +2;g(x)= sqrt x-1

OpenStudy (anonymous):

x=13

OpenStudy (kropot72):

To solve the logarithmic equation write it as log(6x + 2) = log 10 + log(x - 5) ..............(1) This can be done because 1 = log to base 10 of 10 Simplifying (1) gives log(6x + 2) = log 10(x - 5) Can you solve for x now?

OpenStudy (amoodarya):

yes bases are the same so 6x+2=10(x-5) simplify x=13